OpenStudy (cicilybailey14):

Help me!! Suppose you are a math teacher and you have just taught a lesson on exponents. A student asks you to review some worked problems. Look over the problems and correct any mistakes. Explain the errors in the student's reasoning. (4a^2b^5)/(2a^6b^2) = 4/2 * a^2/a^6 * b^5/b^2 = 2 * a^4 * b^3 = 2a^4b^3

OpenStudy (cicilybailey14):

@ribhu ?

OpenStudy (ribhu):

whats the question exactly looks like is unclear to me

OpenStudy (cicilybailey14):

I know what you mean..i'm suppose to fix the equation for the student correct mistakes in the equation if there is any.

OpenStudy (igreen):

I see one mistake.

OpenStudy (igreen):

\(\dfrac{4a^2}{2a^6}\) For this section, we subtract the exponents, giving us 2 - 6, not 6 - 2..you put 4..

OpenStudy (igreen):

\(2a^\color{red}4b^3\)

OpenStudy (cicilybailey14):

I didn't put 4 my teacher did lol but thanks!!!

OpenStudy (igreen):

Oh..

OpenStudy (igreen):

Nevermind..lol.

OpenStudy (cicilybailey14):

So thats not the answer...im suppose to know what the fake "Student" didi wrong was that it?

OpenStudy (igreen):

We divide all the whole numbers together, and subtract all the exponents with like variables. \(\dfrac{4a^2b^5}{2a^6b^2} \rightarrow \dfrac{4}{2} a^{2-6}b^{5-3} \rightarrow 2a^{-2}{b^2}\)

OpenStudy (igreen):

Sorry that should be \(2a^{-4}b^2\)

OpenStudy (igreen):

Yeah, he subtracte 6 - 2, when he should've subtracted 2 - 6.

OpenStudy (igreen):

*subtracted